You are using a wheel and axle as a simple machine. You applied a force of 30 lb to the rim 

of the wheel so that it can lift a load of 240 lb. The diameters of the wheel and axle are 3 feet 

and 4 inches, respectively. Determine the efficiency of the machine


1
Expert's answer
2021-03-09T15:30:20-0500

"\eta F_1d_1=F_2d_2,"

"\eta =\frac{F_2}{F_1}\cdot \frac{d_2}{d_1},"

"\eta =\frac{240}{30}\cdot \frac{4}{12\cdot 3}=0.889=88.9\%."